On an RC circuit where:

the units of E (t) are Volts. Find the charge in Colulombs on the Capacitor at time t = 9 seconds. Take q (0) = 0.1 4.- En un circuito RC donde Cr. 400 F, R=600 N2 , y Elt) 4 si 0 <t < 3, -4 si 3 <t < 6, 0 si 6 <t las unidades de E(t) son
Answer #1

R=6000 Elt) estof 400 Give 91050 Vlo) = 0 Elt): 4 V; OF oc tc3 -4V; 3 ste 6 tz 6 Elt) 41 3 6 -4V Case is for out < 3 sec V.(tLt 3 7 e Volt ) = Vreo) ( ) + V6 (21) (l-ett-3 6 1100) = -4 & TE 3 2 Volt) -4 (2-24) + 3.46 € 14:3) :-) Vc (6) = -4/4-2 6*??Dear student I hope it will help you. Please don't forget to give thumbs up. Thank you